What is the mole fraction of HCl in an aqueous solution that contains 44.2g HCl and 88.6g H20?


1
Expert's answer
2021-03-03T01:57:47-0500

Moles of HCL "x_{HCL}=""\dfrac{\text{Mass of HCL}}{\text{Molar Mass}}=\dfrac{44.2}{36.5}=1.21"



Moles of "H_2O \text{ }x_{H_2O}=\dfrac{\text{Mass of } H_2O}{\text{Molar mass}}=\dfrac{88.6}{18}=4.92"



Mole fraction of HCL "=\dfrac{x_{HCL}}{x_{HCL}+x_{H_2O}}=\dfrac{1.21}{1.21+4.92}=\dfrac{1.21}{6.132}=0.1973"